if 10 be added to 4 times a certain number , the result is 5 less than five times the no.. find the no.
Answer by ankor@dixie-net.com(12704)

Let x = "a certain number"
:
Write an equation for what it says.
"10 be added to 4 times a certain number, the result is 5 less than five times the no."
4x + 10 = 5x - 5
:
10 + 5 = 5x - 4x
:
15 = x
;
:
Check solution in the given statement
10 + 4(15) = -5 + 5(15)